Mindy has 49 nickels and dimes altogether worth $2.80. Find out how many she has of each coin
Answer by jorel1380(3719)

n+d=49
5n+10d=280
5n+5d=245
5d=35
d=7
n=42
Mindy has 42 nickels and 7 dimes..